Science News

Compiled from many sources, especially http://www.sciencenews.org/

Headlines . Are there radio signals from Universe’s first stars? . Body clocks and biorhythms . Wild elephants live longer . Mood-boosting drugs make unhappy fish . Love song of the dengue vector mosquito . How much oxygen do climbers need? . Year of Astronomy starts on Jan 15-16

Some of these are discussed in more detail below. Are there radio signals from Universe’s first stars? Our Universe is known to be a little more than 13 billion years old (1 billion equals 1,000 million). This

is when the Universe was formed, from the Big Bang.

Most astronomers believe that the first stars ever to grace the cosmos with light were brutish monsters, hundreds of times more massive than the Sun. They were very bright when they lived, but exploded or collapsed into massive black holes less than a billion years after the Big Bang, never to be seen again.

But they might have left something behind, a buzz of radio waves emitted by high-energy particles just moments before the collapse.

Has that buzz, a cry from the vanished ancestors of our Sun, now been heard?

In 2006, scientists accidentally recorded a mysterious radio static using a set of sensitive radio receivers called Arcade floated 34 km high in the sky on a balloon.

Microwaves are certain kinds of waves that permeate the entire cosmos. At certain frequencies, there is an excess signal, that seems to be left over from the time of the Big Bang itself. This excess is six times larger than all the signals from all galaxies (known and unknown) and so something is pumping a huge amount of energy into the Universe.

The last two years were spent in making sure that this excess could not be explained by conventional signals, for instance, from distant galaxies.

To prevent heat from the Earth’s atmosphere or anything else from contaminating the delicate measurements, the entire instrument array sits in what is called a flying cold tub. This is literally a giant bucket, open at the top and filled with superfluid liquid helium, which cools the antennas to the same temperature as the universe, 2.7 K (about -270 degrees C).

The frequencies they study are hard to study from the ground-based observatories. The excess that is seen looks like radiation from a radio-galaxy, which happens when particles spiral into a magnetic field.

But if it were just galaxies, scientists should also have seen lots of excess heat in the infra-red frequencies, coming from dust clouds, but this was not seen. Scientists are now trying to improve their theoretical understanding of how much microwave (radio) energy is emitted compared to infra-red energy from galaxies, supernova, black holes or other heavenly objects.

If just radio galaxies is not the right explanation, that leaves open the possibility that the signals are coming from the early stars in the Universe, from a time before the Universe produced any dust (which comes as more and more stars collapse and explode).

So the matter is still not settled. But a lot of people are trying to work out the right answer!

Body clocks and biorhythms

The human body is regulated by several internal clocks, which control sleeping and eating patterns among other things.

Try this: For an entire day, forget about the clock. Eat when you?re hungry and sleep when you?re tired. What do you think will happen?

You may be surprised to find that your day is much like most other days. You’ll probably get hungry when you normally eat and tired when you normally sleep. Even though you don?t know what time it is, your body does.

These patterns of daily life are called circadian rhythms, and they are more than just habits. Inside our bodies are several clock-like systems that follow a roughly 24-hour cycle. Throughout the day and night, our internal clocks direct changes in temperature, body chemicals, hunger, sleepiness and more.

These rhythms are unique, that is, different for different people, which is why you might like to stay up late while your sister always wants to go to bed early. But overall, everyone is programmed to feel tired at night and alert during the day.

Scientists have known for a long time that the light of day and the dark of night play important roles in setting our internal clocks. Now, new discoveries are giving scientists insights into how these clocks work.

In fact, it is known that problems arise when we disobey or disregard our normal bio-rhythms. For instance, regularly staying up late can make children do worse on tests and quizzes. Working night-shifts regularly leads to higher rates of heart disease, diabetes and obesity.

Learning about our body clocks may help scientists understand why problems arise when we act out of step with our circadian rhythms.

One way to learn about how our body clocks tick is to mess them up and see what happens. That’s what happened in a recent study.

The researchers brought 10 people to their lab at Harvard Medical School in Boston. The lab was sort of like a timeless chamber. Rooms were dimly lit. There were no windows and no clocks. It was impossible to know what time it was.

Participants were allowed to sleep only when the scientists said it was OK. The study subjects ate only at designated mealtimes. They were given a precisely calculated number of calories, designed to meet their needs. And they had to finish everything on their plates.

The experiment lasted for 10 days. Participants didn’t know the design of the experiment. In particular, they didn’t know that they were living a 28-hour day instead of the usual 24. With that unusual schedule, they ended up eating and sleeping at all different times of day – and different times of the body clock – over the course of the study.

The most interesting result of the study involved a hormone called leptin. Hormones are the body’s messenger molecules. Leptin, in particular, sends a fullness message to the brain. As you eat, leptin levels rise until you feel like you’ve eaten enough.

When people in the study slept during the day and ate at night, however, leptin levels dropped. That suggests that people who follow unusual schedules are less likely to feel full after eating.

If given unlimited amounts of food, these people would probably eat more and crave more junk food, the researchers predict. As a result, they could gain weight and develop weight-related health problems, such as diabetes and heart disease. Other studies support that prediction.

Other studies on mice show that the food clock and the master clock reside in different part of the brain. Figuring out where the food clock is, especially in humans, will help scientists better understand how it works.

Wild elephants live longer

Scientists have known that elephants in zoos often suffer from poor health. They develop diseases, joint problems and behaviour changes. Sometimes, they even become infertile, or unable to have babies.

To learn more about how captivity affects elephants, a team of international scientists compared the life spans of female elephants born in zoos with female elephants living outdoors in their native lands. Zoos keep detailed records of all the animals in their care, documenting factors such as birth dates, illnesses, weight and death.

These records made it possible for the researchers to analyze 40 years of data on 800 African and Asian elephants in zoos across Europe. The scientists compared the life spans of the zoo-born elephants with the life spans of thousands of female wild elephants in Africa and Asian elephants that work in logging camps, over approximately the same time period.

The team found that female African elephants born in zoos lived an average of 16.9 years. Their wild counterparts who died of natural causes lived an average of 56 years — more than three times as long. Female Asian elephants followed a similar pattern. In zoos, they lived 18.9 years, while those in the logging camps lived 41.7 years.

Scientists don’t yet know why wild elephants seem to fare so much better than their zoo-raised counterparts. Wild elephants live together in large herds and family groups while zoo elephants live alone. That causes a lot of stress. And zoos are never as spacious, so they don’t get as much exercise. Both stress and lack of exercise may cause them to die much earlier than in the wild.

Another finding from the study showed that Asian elephants born in zoos were more likely to die early than Asian elephants captured in the wild and brought to zoos. Mason suggests stress in the mothers in zoos might cause them to have babies that are less likely to survive.

The study raises many questions about acquiring more elephants to keep in zoos. While some threatened and endangered species living in zoos reproduce successfully and maintain healthy populations, that doesn’t appear to be the case with elephants.

Mood-boosting drugs make unhappy fish

When you’re sick, you might take medications to help you fight off infection, lower a fever or clear a stuffy nose. But once those drugs leave your body, they can often find their way into nearby lakes, ponds, rivers and streams.

Drugs end up in a body of water because you excrete them in urine. When you flush a toilet, the wastewater travels to a treatment plant. There, bacteria and other material are filtered out and the cleaned water is returned to natural bodies of water. The trouble is, wastewater treatment plants don’t filter out drugs. Some people even flush unused drugs down the toilet, only adding to the problem.

While medications are meant to help a person feel better, they’re not good for wildlife. Over the past several years, scientists have begun to test how common drugs are in freshwater ecosystems. Researchers also are starting to learn more about how medications meant for humans affect the animals that accidentally ingest the drugs.

Recently, several scientists tested how a group of drugs called antidepressants affects freshwater fish. For many people with an illness called depression, antidepressants can be lifesavers. People with depression may feel sad or anxious for extremely long periods of time, lose interest in activities they once enjoyed and have difficulty sleeping or concentrating. Antidepressants help improve these symptoms for some people.

Several years ago, researchers discovered that some species of fish living near wastewater treatment plants had antidepressants in their brains.

The researchers found that antidepressants affect fish species in numerous ways, from diminishing their response to predators to slowing down their prey-hunting techniques.

Antibiotics, anti-inflammatories, and even caffeine all make their way through water treatment plants back into the environment. What happens to fish and other animals when they’re exposed to all of these drugs in combination? For now, nobody knows.

Love song of the dengue vector mosquito

Mosquitoes use their own kind of song to find a compatible mate. New research shows that male and female mosquitoes sing duets of matching love songs by vibrating their wings. The annoying sounds give researchers some interesting new ways to think about courtship behaviour in insects.

The study focuses on male and female Aedes aegypti mosquitoes, which are carriers of dengue and yellow fever. They are found to change the pitch of their buzzing to match each other’s harmonics.

(You may have noticed that adult males and females have voices at different pitch (frequency). When a man talks at a higher pitch, it is still difficult to match a woman’s pitch. It is easier to match not the same frequency but double (or any multiple) of that frequency. These are called harmonics.)

A female mosquito’s buzz, produced by vibrating her wings at a certain rate, is irresistible to males. Scientists have long thought that male mosquitoes could hear just enough sound to locate a female, while females were thought to be totally deaf. Now it has been found that both can hear.

In the new experiments, researchers delicately tethered live mosquitoes to the ends of flexible wires, and recorded the tones made by the wings as a male and female mosquito came within a few centimeters of each other.

The fundamental tones for each mosquito didn’t change very much; females still produced a fundamental 400-hertz tone and males a 600-hertz tone. But each mosquito produced a faint harmonic note, right around 1200 hertz (twice the base note for a male and thrice that of the female), that were closely in sync with each other.

http://www.sciencenews.org/view/generic/id/39763

How much oxygen do climbers need?

Avalanches, strong winds and sub-zero temperatures aren’t the only extremes endured by those who climb Mount Everest. Hypoxia, a lack of oxygen that can lead to cell death, also threatens climbers.

But a study of people ascending Mount Everest’s slopes suggests that some humans are especially tolerant of low oxygen levels, perhaps because their bodies use oxygen more efficiently. These findings could help in the treatment of critically ill patients struggling to breathe in hospitals.

The blood oxygen levels of the four tested climbers near the peak at 8,400 ms (the peak is at 8,848 ms) were startlingly low, the lowest a mere 19.1 millimeters of mercury, the researchers report. In patients, levels below 60 mm are cause for concern!